OpenStudy (anonymous):

Write the following equation in the form y=mx+b. 1. 8x-2y=6 2. -y+2x-7=0

OpenStudy (anonymous):

Okay so you want to make y the subject. First thing you do is moving all the things that don't affect y to the other side of the = sign. Here the order of operations is handy, as you should deal with those who will affect y last. Example: in 1. y will be multiplied by -2 before being subtracted from 8x, so you move 8x first. -2y = 6 - 8x 2y = -6 + 8x where you simply multiply both sides by - y = -3 + 4x 2. -y + 2x - 7 = 0 -y = 7 - 2x after adding seven to both sides and subtracting 2x from both sides y = 2x - 7 again multiplying by the negative: -